OpenStudy (anonymous):

A cyclist rode the first 28 mile portion of his workout at a constant speed. For the 20 mile cooldown portion of his owrkout, he reduced his speed by 2 miles per hour. Each portion of the workout took the same time. Find the cyclist's speed during the first portio and find his peed during the cooldown portion.

OpenStudy (kendricklamar2014):

s = 20 mph, his initial speed then 20 - 2 = 18 mph his cool-down speed

OpenStudy (dan815):

samething as before 28/x = 20/(x-2)